OpenStudy (anonymous):

1/x = 2+1/y make x as a subject

OpenStudy (zarkon):

reciprocate both sides

OpenStudy (anonymous):

Zarkon's right, reciprocating will get you the right answer. Do you know what "reciprocating" is?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

does making x as a subject means x = (whatever the equation is?)

OpenStudy (anonymous):

\[1/x = 2+ (1/y) \] SOLVE for the terms in the right side of the equation \[1/x=(2y+1)/y\] Cross multiply \[y= 2xy+x\] Then factor y= x(2y+1) Since we're looking for x \[x= y/(2y+1)\]

OpenStudy (anonymous):

Doing the reciprocal is taking the fraction and flipping it. Cross multiplying is also a good way, as herschelle pointed out. Usually much easier, actually.

OpenStudy (anonymous):

thank you